OpenStudy (anonymous):

f(x)=log2 (x)^2/(x-1) and its bases other than e. i am really confused with this one, (x-2)/(ln2)(x)(x-1) is the answer and every time i find my answer its not right please help

zepdrix (zepdrix):

This is really confusing, what is the question?

zepdrix (zepdrix):

\[\Large\rm f(x)=\log_2\left(\frac{x^2}{x-1}\right)\]Is this the function..?

zepdrix (zepdrix):

And what are you trying to do? Derivative? Need more information.

OpenStudy (anonymous):

yes, it was a mess to type sorry

OpenStudy (anonymous):

find the derivative of the function, and with ln

zepdrix (zepdrix):

Apply change of base before taking a derivative:\[\Large\rm \log_2(a)=\frac{\ln(a)}{\ln(2)}\]

zepdrix (zepdrix):

\[\Large\rm f(x)=\frac{\ln\left(\frac{x^2}{x-1}\right)}{\ln(2)}\]Understand what I did there? :o

OpenStudy (anonymous):

i see the ln 2 but you would put the function on the top?

zepdrix (zepdrix):

\[\Large\rm \log_b(\color{royalblue}{c})=\frac{\ln(\color{royalblue}{c})}{\ln(b)}\] Applying it gives us:\[\Large\rm \log_2\left(\color{royalblue}{\frac{x^2}{x-1}}\right)=\frac{\ln\left(\color{royalblue}{\frac{x^2}{x-1}}\right)}{\ln(2)}\]

zepdrix (zepdrix):

I'm not exactly sure what you were asking. Hopefully the color coding helps a little bit.

zepdrix (zepdrix):

Yes, this is all pre-work before we take a derivative.

zepdrix (zepdrix):

Now pull the denominator out of the way since it's just a constant,\[\Large\rm f(x)=\frac{1}{\ln 2}\cdot\ln\left(\frac{x^2}{x-1}\right)\]

zepdrix (zepdrix):

So from here, take the derivative the way you remember. Apply your rule for derivative of natural log, then chain rule into quotient rule.

zepdrix (zepdrix):

No good? =O Too confusing?

OpenStudy (anonymous):

well i see how to use the quotient rule but the chain rule?

zepdrix (zepdrix):

Yah maybe that was sloppy the way I said it. Chain rule is just telling you that you need to multiply by the derivative of the inner function. So chain rule is saying -> Hey after you take derivative of the natural log, apply quotient rule to the inside. Maybe I shouldn't had listed it like a "step". It's just a command telling you to apply quotient rule next.

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{1}{\left(\frac{x^2}{x-1}\right)}\color{royalblue}{\left(\frac{x^2}{x-1}\right)'}\]Chain rule is what makes this blue part show up^ The way you actually DEAL WITH IT, is quotient rule, as you indicated.

OpenStudy (anonymous):

i get with the quotient rule \[x^{2}-1/(x-1)^{2} that correct?\]

zepdrix (zepdrix):

For just the quotient rule portion? Hmm I'm coming up with:\[\Large\rm \frac{x^2-2x}{(x-1)^2}\]Hmm let's see what went wrong... This will let me double check my own work as well.

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left(\frac{x^2}{x-1}\right)'}=\frac{\color{royalblue}{(x^2)'}(x-1)-x^2\color{royalblue}{(x-1)'}}{(x-1)^2}\] \[\Large\rm =\frac{\color{orangered}{(2x)}(x-1)-x^2\color{orangered}{(1)}}{(x-1)^2}\]

zepdrix (zepdrix):

\[\Large\rm =\frac{2x^2-2x-x^2}{(x-1)^2}\]

zepdrix (zepdrix):

I think I did that correctly. Did you forget to distribute the 2x maybe?

OpenStudy (anonymous):

yeah, your right

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{1}{\left(\frac{x^2}{x-1}\right)}\cdot\color{royalblue}{\left(\frac{x^2}{x-1}\right)'}\]So I guess we end up with:\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{1}{\left(\frac{x^2}{x-1}\right)}\cdot\color{orangered}{\frac{2x^2-2x-x^2}{(x-1)^2}}\]

zepdrix (zepdrix):

Which can be simplified a little bit further.

OpenStudy (anonymous):

yeah i was just abou to say 2x^2-x^2 can be X^2

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{1}{\left(\frac{x^2}{x-1}\right)}\cdot\frac{x^2-2x}{(x-1)^2}\] How bout that middle part? Remember what to do when you divide by a fraction?

OpenStudy (anonymous):

do the reciprocal

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{x-1}{x^2}\cdot\frac{x^2-2x}{(x-1)^2}\]Ok good. That should make canceling things out much easier.

zepdrix (zepdrix):

In the last piece, factor an x out of each term in the numerator,\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{x-1}{x^2}\cdot\frac{x(x-2)}{(x-1)^2}\]You should see the dominoes all tipper over at this point c:

zepdrix (zepdrix):

See some stuff you can cross out?

OpenStudy (anonymous):

(x-1) correct?

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{\cancel{x-1}}{x^2}\cdot\frac{x(x-2)}{(x-1)^{\cancel{2}1}}\]Yah, there's something else though.

OpenStudy (anonymous):

the x's too

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\ln 2}\cdot\frac{1}{x^{\cancel{2}1}}\cdot\frac{\cancel{x}(x-2)}{(x-1)}\]Ah yes good.

zepdrix (zepdrix):

Now just combine all the pieces and it should match the answer which was given.

zepdrix (zepdrix):

Kind of a tough problem. No only do you have a few derivative steps to work through, but you also had to remember your log rule, and then a ton of algebra steps to simplify it down.

OpenStudy (anonymous):

yeah, thanks zepdrix, that help :)

zepdrix (zepdrix):

cool c: